OpenStudy (anonymous):

Suppose a bag of marbles has 4 green, 2 red, 5 yellow, 1 brown, and 7 blue marbles. What is the probability of picking a green marble, replacing it, and then picking a brown marble?

OpenStudy (anonymous):

let's start with the sample space. S= all possible outcomes what is n(S) = ?

OpenStudy (anonymous):

s=19 then? so then would it be 4/19?

OpenStudy (anonymous):

nope.. next we have to set the events that they are looking for. A= picking the green marble as first pick n(A) = ?

OpenStudy (anonymous):

don't ignore the second event that they mention: "then picking a brown marble" so event B= picking a brown marble n(B)=?

OpenStudy (anonymous):

since these events are independent.. the formula that you have to use is..\(P(A\ and\ B)= P(A) \times P(B)\) \(P(A)=\frac{n(A)}{n(S)}\) and \(P(B)=\frac{n(B)}{n(S)}\)